For this assignment, you will continue researching your chosen company. Compose a paper on the financials of your chosen firm; you will choose two ratios from each category (liquidity, profitability, and solvency) and conduct an analysis as well as highli

My company is AMAZON


For this assignment, you will continue researching your chosen company. Compose a paper on the financials of your chosen firm; you will choose two ratios from each category (liquidity, profitability, and solvency) and conduct an analysis as well as highlight key areas/trends of the income statement and balance sheet. Be sure to answer the following questions in your response:

  • What do the ratios reflect?

  • Does the company appear healthy and headed in the right direction? Explain.

  • How does the company compare to another firm in the same industry in regards to financial metrics (for instance, a comparison of Target versus Walmart based on the basic ratio categories mentioned above)?

Your paper must be at least two pages in length, and you must use at least two sources. Adhere to APA Style when creating citations and references for this assignment.
Answer #1

LIQUIDITY RATIOS

Liquidity ratios are used to measure your company's ability to cover its expenses. The two most common liquidity ratios are -

Current ratio & Quick ratio.

Both are based on balance sheet items.

Current Ratio

The current ratio is calculated as Total Current assets /Total Current liabilities

For AMAZON, the current ratio can be calculated from their balance sheet items.

Total current assets - $96334 million

Total current liabilities- $ 87812 million

Current ratio - 96334 / 87812 = 1.097 i.e. 1.10 (AMAZON)

A common rule of thumb is that a "good" current ratio is 2 to 1. This current ratio indicates that the company's cash is optimally utilized.

Quick Ratio

Quick ratio= Quick assets / Current Liabilities

"Quick" assets include cash, stocks and bonds, and accounts receivable (in other words, all of the current assets on the balance sheet except the inventory)

Quick assets - Current Assets - inventory

For AMAZON

Total current assets - $96334 million

Inventory - $20497 million

Total current liabilities- $ 87812 million

QUICK RATIO =(96334- 20497)/87812 = 0.86

Quick ratios between 0.5 and 1 are considered to be satisfactory—as long as the collection of receivables is not expected to slow down. Therefore Amazon's quick ratio is adequate.

PROFITABILITY RATIOS

Profitability ratios are used to measure a company’s ability to earn a profit relative to its sales revenue, operating costs, balance sheet assets and shareholders’ equity. These financial metrics show how well companies use their existing assets to generate profit and value for its owners and shareholders.

Two main profitability ratios are-

Return On Assets and Return on Equity

RETURN ON ASSETS

Return on asset ratio or Return on Investment ratio tells about the efficiency. It indicates how good is your company in turning its investments into profit.

It can be calculated as-

Net income/Total assets

For amazon as per its income statement and balance sheet-

Net income =$11588 million

Total assets=$ 225248 million

ROA (11588/225248)*100= 5.14

The ROA of 5.14 is considered as good and indicates that the company can earn income efficiently from its available assets.

RETURN ON EQUITY

ROE shows how well a company can use its shareholders investments to generate profits. It is a measure of a shareholder’s return on their investment.

ROE can be calculated as-

(Net income/ Average Shareholders' Equity)*100

As per income statements and balance sheet of Amazon

Average Shareholders' Equity =$62060 million

Net income =$11588 million

ROE = (11588/62060)*100 = 18.67

The ROE of 18.67is quite high and indicates that the management is doing a good job at using its investors’ funds to generate return.

SOLVENCY RATIOS

Solvency ratios measure the stability of the company and its ability to repay debt.It also give a strong indication of the financial health and viability of the company.

Two main solvency ratios are-

Debt to Equity Ratio and Debt ratio

Debt To Equity Ratio

The debt to equity ratio measures the relationship between the long-term debt of a company and its total equity

It can be calculated as-

D/E ratio= Long term debt / total share holder funds

Lond Term Debt = Debentures + Long Term Loans

Shareholders Funds = Equity Share Capital + Preference Share Capital + Reserves – Fictitious Assets

D/E ratio for Amazon is-

Long term debt - $23.414million

Shareholder funds - $ 62060 million

D/E = 0.3773

Ideal D/E ratio is 2:1 that is twice the amount of debt to equity.But it cannot be said if its wrong to maintain a low D/E ratio as low ratio means the firm is more financially secure and equity is diluted.

DEBT ratio

This ratio measures the long-term debt of a firm in comparison to its total capital employed or net fixed assets.It shows the financial leverage of the firm

Debt ratio for amazon- Long term debt/ net assets

Long term debt-$23.414million

Net fixed assets- $87495

Debt ratio= 0.267

The low ratio of 0.267 points to a more financially stable business.

COMPARISON OF AMAZON TO WALMART

Liquidity ratio analysis-

On this ground Amazon has a current ratio of 1.10 whereas Walmart has a current ratio of 0.79 which is lower than that of Amazon and mostly higher current ratio is preferred to be good. Thus, Amazon leads in this analysis.

Profitability Ratio Analysis-

Here key ratio is considered to be ROE. So Amazon's ROE is 18.67% whereas Walmart's ROE is precariously 7.2%.

And usually higher ROE is considered to be the better one.So ,here also Amazon takes the lead over Walmart.

Solvency Ratio Analysis-

Here key ratio is considered to be D/E ratio. So Amazon's D/E is 37.73% whereas Walmart's D/E is 97.01%.Any ratio under 100% is considered to be better as higher indicates that its debt load has overtaken the value of its equity.

Here also Amazon has the lower D/E which is preferred in assessing the financial security of the company.

Thus , as per the ratios mentioned above Amazon is considered to be healthier and is headed in the right direction.Because all the ratios are as per the ideal scenarios prescribed for it.


answered by: anonymous